The first chemical step in the extraction of silver from silver sulfide containing ores involves moving the
silver from the insoluble silver sulfide within the ore, Ag 2S, into a soluble silver salt by complexing with
sodium cyanide:

Ag 2S(s) + 4 NaCN(aq) → 2 NaAg(CN) 2(aq) + Na 2S(aq)


A 1.14­kg sample of ore containing Ag 2S is treated with NaCN. The treatment requires 785 mL of
0.0424 M NaCN to extract all the silver from the Ag 2S contained in the 1.14­kg ore sample. Find the
mass percent of silver in the 1.14­kg ore sample.
Answer #1

The mass of Ag is calculated:

m Ag = 0.0424 M * 0.785 L * (2 mol Ag / 4 mol NaCN) * (108 g Ag / 1 mol) = 1,797 g

The percentage is calculated:

% Ag = g Ag * 100 / g ore = 1,797 * 100/1140 = 0.16%

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.